OpenStudy (hitaro9):

Show that the sequence {(4n-1)/(n-2)} is bounded for n>3

OpenStudy (hitaro9):

I don't think this is too tough I'm just not seeing it though

OpenStudy (phi):

divide top and bottom by n

OpenStudy (phi):

you should get \[ \frac{ 4 - \frac{1}{n}}{1-\frac{2}{n}}\]

OpenStudy (phi):

and 1/n for large n becomes very small. in the limit as n approaches infinity , 1/n approaches 0

OpenStudy (hitaro9):

Right that makes sense, thank you.